Grade 9Mechanics

a window is 50 cm long. a stone is falling from a height of 40 cm above the window . it crosses the window in ??
the answer is 3/7 secs but how ???

DEEPNIL PATHAK
9 Years agoGrade 9

2 Answers

Sorabh Gulia
8 Years ago
Hlo friend try this
Use the kinematic equations as follows:
SUVAT-- s-displacement, u-initial velocity, v-final velocity,a- acceleration, t-timev=u + atv2 = u2 + 2ass=1/2(u+v)ts=ut + 1/2at2
The values for this questions are as follows:
s= 40cm and 50cm
a= 9.8m/s2 (due to gravity)
u=0m/s
The time it takes to reach the top of the window:
s=0.4m
s=ut+1/2at^2
0.4=0xt + 1/2 x 9.8 x t^2
0.4=4.9t^2 
t^2= 0.4/4.9
= 0.08163
t= 0.285 s                      
The time taken to cross the window:
s=ut+1/2at^2
0.5= (0 x t) + 0.5x9.8xt^2
4.9t^2=0.5 
t2=0.5/4.9  
t2= 0.1020
t = 0.1010 seconds
The stone crosses the window in 0.1010 seconds
Utkarsh Khichariya
8 Years ago
Given : 
S= 40 cm = 0.4 m
S2 = 50 cm = 0.5 m
S=S1+S2 = 0.4+0.5 = 0.9 m = 9/10 m
a = 9.8 m/s2
uinitial = 0 
Therefore from 2nd equation of motion ;
s= ut+ ½ ×at2
9/10 = 0×t + ½ × 9.8 × t2 
9/10 = 0+4.9×t2
t = 9/49
t = 3/7 seconds